Convert 41.9degrees and 74.28degrees to the nearest degree, minutes and seconds. Please show working
Found 2 solutions by Theo, Alan3354:
Answer by Theo(13342)

as far as i know, there are 60 minutes in 1 degrees and 60 seconds in 1 minute.

41.9 degrees is equal to 41 degrees plus .9 degrees.
.9 degrees is equal to .9 * 60 = 54 minutes.
therefore 41.9 degrees is equal to 41 degrees 54 minutes.

74.28 degrees is equal to 74 degrees plua .28 degrees.
.28 degrees is equal to .28 * 60 = 16.8 minutes.
16.8 minutes is equal to 16 minutes plus .8 minutes
.8 minutes is equal to .8 * 60 = 48 seconds.
therefore 74.28 degrees is equal to 74 degrees 16 minutes 48 seconds

Answer by Alan3354(69443)

Convert 41.9degrees and 74.28degrees to the nearest degree, minutes and seconds
---------------------
0.28 degs = 0.28*60' = 16.8'
0.8' = 0.8*60" = 48"
---> 74degs 16' 48"